Sea-Salt Restaurant borrows R100000

 for five years from a bank to upgrade its premises. The bank charges interest at 12%

 per annum compounded monthly over five years. What is the equal monthly repayments of Sea-Salt Restaurant to the bank?


1
Expert's answer
2022-02-17T13:18:21-0500

The formula for finding the monthly loan payment is as follows:

"A=S\times(i+\frac{i}{(1+i)^n-1})"

where S is the loan amount

S=100 000

n - number of months of loan repayments:

i - monthly loan rate

"i=\frac{12}{12}=1"


"A=100 000\times(0.01+\frac{0.01}{(1+0.01)^{60}-1})=2224.45"